The lifetime of a machine is continuous on the interval (0, 40) with probability density
function f, where f(t) is proportional to (t+ 10)−2
, and t is the lifetime in years. Find
the proportionality constant that makes the f(t) a probability function.
1
Expert's answer
2020-06-01T16:11:49-0400

By the definition, "f(t)=(t+10)-2" is the probability density function if



"\int\limits_{-\infty}^{+\infty}f(t)dt=1\to \int\limits_0^{40}\left(t+10-2\right)dt=\\[0.3cm]
=\int\limits_0^{40}\left(t+8\right)dt=C\cdot\left.\left(\frac{t^2}{2}+8t\right)\right|_0^{40}=\\[0.3cm]
C\cdot\left(\frac{1600}{2}+320\right)=1\to C\cdot1120=1\longrightarrow\\[0.3cm]
\boxed{C=\frac{1}{1120}}"

Comment: Perhaps the probability distribution function was incorrectly indicated and actually meant "f(t)=1\/(t+10)^2" , then



"\int\limits_{-\infty}^{+\infty}f(t)dt=1\to \int\limits_0^{40}\left(\frac{1}{(t+10)^2}\right)dt=\\[0.3cm]
-C\cdot\left.\left(\frac{1}{t+10}\right)\right|_0^{40}=-C\cdot\left(\frac{1}{50}-\frac{1}{10}\right)=1\\[0.3cm] -C\cdot\frac{1-5}{50}=1\longrightarrow C=\frac{50}{4}=\frac{25}{2}=12.5\\[0.3cm]
\boxed{C=\frac{25}{2}=12.5}"


ANSWER



"\text{If}\,\,\,f(t)=(t+10)-2\longrightarrow C=\frac{1}{1120}\\[0.3cm]
\text{If}\,\,\,f(t)=\frac{1}{(t+10)^2}\longrightarrow C=\frac{25}{2}=12.5\\[0.3cm]"